OpenStudy (anonymous):

A boy standing on a 64 foot cliff drops a coin off the cliff. If the equation of the height h of the coin t seconds after it drops is h = -16t2 + 64, about how many seconds will the coin take to reach the bottom of the cliff? A. 1 second B. 1.5 seconds C. 2 seconds D. 2.5 seconds

OpenStudy (anonymous):

okay what is the height of the bottom of the cliff?

OpenStudy (anonymous):

lets go over this together.

OpenStudy (anonymous):

probably 64!

OpenStudy (anonymous):

try again.

OpenStudy (anonymous):

haha :O ok ya lets go over it.

OpenStudy (anonymous):

you can almost see that 2 squared is 4, and 4 * 16 = 64

OpenStudy (anonymous):

the equation states that h is the height of the coin t seconds after it is dropped. You are asked to find when it reaches the bottom. So what is the height of the bottom?

OpenStudy (anonymous):

the cliff is 64 ft. right? lol cuz u said whats the height of the bottom of the cliff i think

OpenStudy (anonymous):

yes, I know what I asked. So, answer me, what do you think the height of the bottom of the cliff is?

OpenStudy (anonymous):

64!

OpenStudy (anonymous):

if I am 6 feet high, how high is the bottom of my feet from the ground?

OpenStudy (anonymous):

ok uhmm blonde moment rite now how bout 16?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

YEAAH

OpenStudy (anonymous):

haha wow blonde moment

OpenStudy (anonymous):

throw h=0 in and solve for t!

OpenStudy (anonymous):

Yes!. so the moment the coin touches the bottom of the cliff, it is 0 feet high from the ground.

OpenStudy (anonymous):

rite.

OpenStudy (anonymous):

craig is trying to rush through it lol

OpenStudy (anonymous):

so equate h to 0 and solve the equation, that is, 0 = -16t^2+64

OpenStudy (anonymous):

ah! ok one moment keep typing watever ur typing.

OpenStudy (anonymous):

or, 16 t^2 = 64. or t^2 = 4 or t = +2 or -2. Since time cannot be negative, t = +2. So it takes 2 seconds to reach the bottom.

OpenStudy (anonymous):

OH!!!! ok ok

OpenStudy (anonymous):

that makes sense! (: thx!! xD

OpenStudy (anonymous):

okay great! you are welcome.